OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + x2 - 20x ?

OpenStudy (anonymous):

@satellite73 mind helping me quick? i dont know how to do this.....

OpenStudy (anonymous):

\[x^3+x^2-20x=0\] \[x(x^2+x-20)=0\] \[x(x+5)(x-4)=0\] and you should be good to go

OpenStudy (anonymous):

Thanks got it!!